OpenStudy (studystagram):

For what value of k, the system of equations 2/5x+1/ky=4 and 3/5x - 6y =7 will have unique solution?

OpenStudy (anonymous):

\[\dfrac{2}{5} x + \dfrac{1}{k} y = 4\]\[\dfrac{3}{5} x - 6y = 7\]--- Do you know the condition for a unique solution?

OpenStudy (studystagram):

Yup...I was just confused as to what to do with the 1/ky thing. Should I leave it like that or do i have to simplify or something?

OpenStudy (anonymous):

You will have to use that condition. Can you recall what it is?

OpenStudy (studystagram):

Yeah I know that... 

OpenStudy (anonymous):

Apply the same condition here.

OpenStudy (studystagram):

I want to know the answer, dude

OpenStudy (anonymous):

Then find it.

OpenStudy (studystagram):

That really helps. thanks

OpenStudy (anonymous):

\[\large \dfrac{\frac{2}{5}}{\frac{3}{5}}\ne \dfrac{\frac{1}{k}}{-6}\]\[\Rightarrow \dfrac{2}{3} \ne - \dfrac{1}{6k}\]

OpenStudy (studystagram):

THIS IS EXACTLY WHERE I WAS CONFUSED. THANK YOU FOR SOLVING!!

OpenStudy (anonymous):

Are you being sarcastic?

OpenStudy (studystagram):

no lol

OpenStudy (anonymous):

One thing is that \(k \ne 0\) because you can never divide by zero. And from the above equation, you will find another value that \(k\) cannot hold.

OpenStudy (studystagram):

okay. thanks again!